### v1

Abstract
Non-cooperative interactions commonly occur in multi-agent scenarios such as car racing, where an ego vehicle can choose to overtake the rival, or stay behind it until a safe overtaking “corridor” opens. While an expert human can do well at making such time-sensitive decisions, the development of autonomous agents capable of rapidly reasoning about complex, potentially conflicting options is yet to be fully addressed. The recently developed nonlinear opinion dynamics (NOD) model shows promise in enabling fast (i.e., at an exponential rate) opinion formation and avoiding safety-critical deadlocks. However, it remains an open challenge to determine the model parameters of NOD automatically and adaptively, accounting for the ever-changing environment of interaction. In this work, we propose for the first time a learning-based, game-theoretic approach to synthesize a Neural NOD model from expert demonstrations, given as a dataset containing (possibly incomplete) state and action trajectories of interacting agents. The learned NOD can be used by existing dynamic game solvers to plan decisively while accounting for the predicted change of other agents’ intents, thus enabling situational awareness in planning. We demonstrate Neural NOD’s ability to make fast and robust decisions in a simulated autonomous racing example, leading to tangible improvements in safety and overtaking performance over state-of-the-art data-driven game-theoretic planning methods.
Keywords:
nonlinear opinion dynamics inverse games neural synthesis
1 Introduction
As autonomous multi-agent systems evolve toward an unprecedented complexity, especially in high-stakes scenarios such as car racing, the necessity for reliable decision-making in real-time becomes paramount. Rapid decision-making is critical in these settings, not only for competitiveness but also for ensuring safety during complex, close-proximity interactions. For example, in autonomous car racing, the ego vehicle must decide, in a split second, between overtaking a rival car without crashing into the rival or strategically trailing it to capitalize on emerging opportunities. While existing game-theoretic approaches [5] effectively capture non-cooperative interactions, they often fall short in real-time calibration of strategies to account for the evolving interactions with other agents. Although multi-agent reinforcement learning (RL) and imitation learning approaches, such as [47, 53], show promise in performance and scalability, they may require massive amounts of data to mitigate generalization issues [43], and it remains unclear how these methods can ensure fast adaptation to changing environments and avoid safety-critical indecisions due to lack of interpretability in decision-making.
Figure 1: Rapid and resolute decision-making is essential for non-cooperative multi-agent interactions like car racing. Top: During the 2021 Formula 1 Italian Grand Prix, a fatal collision occurred involving championship contenders Max Verstappen and Lewis Hamilton. Verstappen was deemed predominantly responsible because, despite the overtaking opportunity closing after Hamilton (orange triangle) led him into the corner, he had options to avoid the collision by slowing down or taking the emergency alternative route (green arrows), but he failed to make a timely decision, continuing along the racing line (red arrow) and ultimately leading to an inevitable collision later on. Middle: A similar scenario arises in simulated autonomous racing when the ego car (red) uses an indecisive policy, hesitating between overtaking the rival (silver) from the inside or outside of the corner (as seen in its planned future motions depicted in transparent snapshots), ultimately resulting in a collision. Bottom: The proposed Neural NOD model reasons split-second strategic interactions between the agents, rendering safe and decisive overtaking maneuvers.
The recently developed NOD model [6] provides a principled framework to describe opinion exchanges among agents and guarantees indecision breaking. More specifically, an opinion state numerically represents agents’ agreement or disagreement on a topic that is relevant to collective decision-making, thus being a natural mathematical tool for modeling multi-robot non-cooperative interactions. NOD achieves rapid opinion formation with nonlinear bifurcation, at which point the process is ultrasensitive to external inputs, e.g., physical interactions among agents. However, it remains an open challenge to select suitable NOD model parameters such that the ego agent can reason how ongoing interactions may drive rapid changes in its own situational awareness and intended course of action as well as those of other agents.
Contributions and Paper Organization. We propose for the first time an inverse-game-based approach to synthesize an NOD model parameterized by a deep neural network (DNN). This model is used for automatic cost tuning by online dynamic game solvers to enable time-sensitive multi-agent decision-making. We provide verifiable conditions under which the model is guaranteed to avoid indecisions. We deploy the Neural NOD learned from real human racing data for simulated autonomous car racing and compare our approach against the state-of-the-art data-driven game-theoretic planning baseline.
The paper is structured as follows. Section 2 discusses related work in multi-agent decision-making. Section 3 reviews dynamic games and NOD. In Section 4, we discuss how to learn a Neural NOD model from data leveraging inverse dynamic games, and use it for split-second decision-making in multi-agent settings. Section 5 presents simulation results in autonomous car racing. We discuss limitations and future work in Section 6. Finally, Section 7 concludes the paper.
2 Related Work
Multi-Agent Motion Planning. Early efforts in multi-agent motion planning focus on developing decentralized or distributed algorithms, leveraging single-agent planning approaches such as graph search [46], model predictive control [12], and trajectory optimization [1]. However, these methods can struggle with general non-cooperative interactions involving self-interested, potentially competing agents. Game-theoretic motion planning explicitly reasons about the coupled interests among agents, and has been widely used for multi-agent non-cooperative interactions, e.g., autonomous driving [44, 13, 50, 24], human–robot interaction [31, 4, 21], and distributed control systems [35, 20, 51]. Computing equilibria is oftentimes computationally intractable for general dynamic games due to their inherent non-convexity involving, e.g., nonlinear system dynamics, non-convex cost functions, and collision-avoidance constraints. To tackle such computation challenges, Fridovich-Keil et al. [14] apply iterative linear-quadratic (ILQ) approximation to solve general-sum dynamic games, enabling real-time computation of approximate Nash equilibria. Recent work [32] extends this approach to compute a mixed Nash strategy that blends task optimality with a multi-modal data-driven reference policy.
In contrast to model-based optimization, data-driven methods have also shown competitive performance in interactive autonomy. GT Sophy [52] demonstrates that a well-trained deep RL policy can win a head-to-head competition against some of the world’s best drivers in Gran Turismo, a popular car racing video game. However, the policy does not explicitly reason the strategic interactions among players. To investigate this gap, Chen et al. [10] combine RL with self-supervised learning that models the opponent behaviors, and show that the resulting policy outperforms methods with manually designed opponent strategies in multi-agent car racing. Song and Lin et al. [47] use curriculum learning to improve the overtaking performance in Gran Turismo races. Recent work [23] leverages belief-space planning to actively reduce opponent uncertainty in adversarial interactions, and uses deep RL to synthesize a best-effort reach-avoid policy. Despite promising performance and scalability, these black-box methods lack game-theoretic interpretations when deployed for complex interactions, and may suffer from generalization issues when data are insufficient [43].
Our proposed motion planning framework combines the best of both worlds by offering a game-theoretically meaningful solution with a model-based solver, which is guided by a Neural NOD model learned from interaction data for fast and decisive motion planning in rivalrous, safety-critical interactions.
Inverse Dynamic Games. Unlike forward games that focus on computing equilibria strategies, inverse dynamic games are useful when information of the game, e.g., cost parameters, is unknown or unspecified, but may be learned from data. Peters et al. [42] leverages differentiable optimization [3] to learn players’ a priori unknown objectives of a non-cooperative dynamic game based on noise-corrupted, partial state observations. The optimization couples past state estimates with future state predictions, thus enabling simultaneous online learning, prediction, and decision-making. Liu et al. [33] further extend this idea to solve constrained inverse dynamic games, which yield generalized equilibria strategies. Another concurrent follow-up work [29] proposes an inverse dynamic game framework assuming all players adopt a feedback Nash equilibrium strategy. However, these inverse game methods mainly focus on learning static cost parameters from a single trajectory. As such, the inverse game needs to be re-solved when the environment (e.g., the initial state) has evolved, posing computational challenges at runtime.
Our approach builds on existing inverse game approaches to learn a Neural NOD model, and use it in forward game solvers to facilitate time-sensitive decision-making in rapidly evolving interactive environments. To the best of our knowledge, our work is the first to apply inverse dynamic games to learn a deep model from many trajectories, which practically generalizes to arbitrary initial states at deployment time.
Nonlinear Opinion Dynamics. NOD is pioneered by Bizyaeva et al. [6] for deadlock-free decision-making in multi-agent interactions. Cathcart et al. [9] use NOD to break social deadlocks in a human–robot corridor passing problem. Amorim et al. [2] propose an NOD-based decision-making framework for robots choosing between different spatial tasks while adapting to environmental conditions, where opinions represent agents’ preference for a particular task. Paine et al. [39] use NOD to make group decisions for autonomous multi-robot systems, demonstrating robust and interpretable collective behaviors in a field test involving multiple unmanned surface vessels. Hu et al. [22] design for the first time a general algorithm to automatically synthesize an NOD model. They propose to construct NOD parameters based on dynamic game value functions, thus making the opinion evolution dependent on agents’ physical states. However, this approach assumes that agents are facing a set of mutually exclusive options, and it cannot effectively integrate data-driven prior knowledge to facilitate online decision-making. For a comprehensive review of NOD, we refer the readers to Leonard et al. [28].
The proposed Neural NOD can be trained over a dataset of diverse interaction trajectories, account for general opinion spaces, and is plug-and-play for any model-based game solver. Compared to previous NOD with static model parameters, the Neural NOD dynamically adjusts its parameters according to the evolving physical states, allowing the ego robot to rapidly adapt its decision to the fast-changing environment.
3 Preliminaries
Notation. We use superscript ${i}$ to indicate agent ${i}\in\mathcal{I}_{a}:=\{1,2,\ldots,{N_{a}}\}$. We define $[N]:=\{1,2,\ldots,N\}$. Given a function $f$, we denote ${\mathbf{J}}(f(\cdot))$ as its Jacobian, and ${\nabla}_{\theta}f$ as the derivative of $f$ with respect to variable $\theta$. Given a matrix $A\in\mathbb{R}^{n\times n}$, let ${}^{+}(\sigma(A))$ be the set of positive real part of eigenvalues of $A$. Define $\operatorname{col}(\cdot)$ as the operator that stacks all the arguments into a column vector.
General-Sum Dynamic Games. We consider an ${N_{a}}$-player finite-horizon, discrete-time dynamic game governed by a nonlinear dynamical system:
${x}_{t+1}={f}({x}_{t},{u}_{t}),$ (1)
where ${x}\in\mathbb{R}^{n_{x}}$ is the joint state of the system, which we assume to be fully observable at runtime, ${u}_{t}:=({u}^{1}_{t},\ldots,{u}^{{N_{a}}}_{t})$ where ${u}^{i}\in\mathbb{R}^{m_{i}}$ is the control of player ${i}$. The objective of each player ${i}$ is to minimize a cost functional:
$J^{i}\left({\pi};\theta^{i}\right):=\textstyle\sum_{k=0}^{T}c_{k}^{i}({x}_{k},{u}_{k};\theta^{i}),$ (2)
where $c^{i}_{k}(\cdot)$ is the stage cost, $\theta^{i}\in\mathbb{R}^{n_{\theta^{i}}}$ is the (possibly unknown) parameter of the stage cost, and ${\pi}:=({\pi}^{1}(\cdot),\ldots,{\pi}^{N_{a}}(\cdot))$ is a tuple containing all players’ control policies, which determines the information structure [5, Ch. 5] of the game, e.g., open-loop if ${u}^{i}_{t}={\pi}^{i}({x}_{0})\in\mathbb{R}^{m_{i}}$ and feedback if ${u}^{i}_{t}={\pi}^{i}({x}_{t})\in\mathbb{R}^{m_{i}}$. Policy tuple ${\pi}$ can constitute different equilibrium types of the game, among which the most common ones are Nash [37] and Stackelberg [48]. In this paper, we do not adhere to a specific information structure or equilibrium type.
Inverse Dynamic Games. When cost parameters $\theta:=(\theta^{1},\ldots,\theta^{N_{a}})$ are initially unknown, we can solve an inverse dynamic game [42, 29, 33] to identify them from trajectory data. This is formulated as a maximum likelihood estimation (MLE) problem:
$\max_{\theta,{\mathbf{x}},{\mathbf{u}}}~{}p(\mathbf{{y}}\mid{\mathbf{x}},{\mathbf{u}}),\quad\text{s.t.}({\mathbf{x}},{\mathbf{u}})\in\Gamma(\theta)$ (3)
where $p(\mathbf{{y}}\mid{\mathbf{x}},{\mathbf{u}})$ is the likelihood of observed trajectory data $\mathbf{{y}}:={y}_{[0:T]}$ given state trajectories ${\mathbf{x}}:={x}_{[0:T]}$ and control sequences ${\mathbf{u}}:={u}_{[0:T-1]}$ that correspond to a (user-specified) equilibrium, and $\Gamma(\theta)$ is the set of all such equilibrium solutions parameterized by $\theta$. MLE (3) can handle corrupted data: the initial state ${x}_{0}$ is not assumed to be known, and the observation data $\mathbf{{y}}$ may be noisy and missing certain entries (e.g., at specific time indices). An inverse game routine such as [42, 29, 33] solves MLE (3) by applying gradient ascent to update the value of $\theta$, where gradient ${\nabla}_{\theta}p(\cdot)$ is computed by differentiating through a forward game solver that produces an equilibrium solution $({\mathbf{x}},{\mathbf{u}})\in\Gamma(\theta)$. Existing inverse game methods predominantly learn a static, context-agnostic cost parameter $\theta$ along a single game trajectory ${\mathbf{x}}$. In Section 4, we extend established inverse game techniques to learn an adaptive, context-aware cost parameter governed by an NOD from a dataset containing diverse gameplay trajectories.
Nonlinear Opinion Dynamics. The NOD model [6] enables fast and flexible multi-agent decision-making. Consider multi-agent system (1), in which each agent ${i}$ is faced with an arbitrarily large (but finite) number of ${N_{o^{i}}}$ options. For every ${i}\in\mathcal{I}_{a}$ and $\ell\in\mathcal{I}_{o^{i}}:=\{1,\ldots,{N_{o^{i}}}\}$, define ${z}^{{i}}_{\ell}\in\mathbb{R}$ to be the opinion state of agent ${i}$ about option $\ell$. The more positive (negative) is ${z}^{{i}}_{\ell}$, the more agent ${i}$ favors (disfavors) option $\ell$. We say agent ${i}$ is neutral (i.e., undecided) about option $\ell$ if ${z}^{{i}}_{\ell}=0$. Compactly, we define ${z}^{i}:=({z}^{i}_{1},\ldots,{z}^{i}_{{N_{o^{i}}}})$, and ${z}:=({z}^{1},\ldots,{z}^{{N_{a}}})$ as the opinion state of agent $i$ and the joint system, respectively. The evolution of opinion state in continuous time is governed by the NOD model, which is an ordinary differential equation:
$\dot{z}^{i}$ $={g}^{i}_{c}({z}^{i})=-D^{{i}}{z}^{{i}}+{b}^{i}+{\lambda}{S}^{i}({z}^{{i}}),$ (4)
where the $\ell$-th entry of the saturation term ${S}^{i}({z}^{i})$ is
${S}^{{i}}_{\ell}(\cdot)={S}_{1}\Big{(}\alpha^{{i}}_{\ell}{z}^{{i}}_{{\ell}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\gamma^{{i}{j}}_{\ell}{z}^{{j}}_{{\ell}}\Big{)}+\sum_{{p}\in\mathcal{I}_{\theta_{i}}\setminus\{{\ell}\}}{S}_{2}\Big{(}\beta_{\ell{p}}^{{i}}{z}^{{i}}_{{p}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\delta^{{i}{j}}_{\ell{p}}{z}^{{j}}_{{p}}\Big{)}$
which satisfies for $r\in\{1,2\}$, ${S}_{r}(0)=0$, ${S}^{\prime}_{r}(0)=1$, ${S}^{\prime\prime}_{r}(0)\neq 0$, and ${S}^{\prime\prime\prime}_{r}(0)\neq 0$. Valid choices for ${S}_{r}(\cdot)$ include the sigmoid function and the hyperbolic tangent function $\tanh$. In NOD model (4), $D^{{i}}=\operatorname{diag}({d}^{i}_{1},\ldots,{d}^{i}_{N_{o^{i}}})$ is the damping matrix with each ${d}^{i}_{\ell}>0$, ${b}^{i}$ represents agent ${i}$’s own bias, ${\lambda}>0$ is the attention value on nonlinear opinion exchange, which is shared across all agents, $\alpha^{{i}}_{\ell}\geq 0$ is the self-reinforcement gain, $\beta^{{i}}_{\ell{p}}\geq 0$ is the same-agent inter-option coupling gain, $\gamma^{{i}{j}}_{\ell}$ is the gain of the same-option inter-agent coupling with another agent ${j}$, and $\delta^{{i}{j}}_{\ell{p}}$ is the gain of the inter-option inter-agent coupling with another agent ${j}$. In order to guide a dynamic game with NOD, we consider the discrete-time variant of the NOD model jointly for all ${i}\in\mathcal{I}_{a}$:
${z}_{t+1}$ $={g}({z}_{t}),$ (5)
which may be obtained by applying time discretization (e.g., forward Euler or Runge-Kutta method) to the continuous-time NOD model (4). While NOD has demonstrated efficacy in multi-agent decision-making, automatic synthesis of its parameters that can change adaptively based on physical states, e.g., $\alpha^{i}_{t}({x}_{t})$, largely remains an open problem. In the next section, we propose a novel inverse-game-based approach that learns a general DNN-parameterized NOD model, which allows the opinion state ${z}$ to be influenced by physical states ${x}$.
Remark 1 (NOD Tunable Sensitivity)
The NOD provides tunable speed and sensitivity, that is, a “fast” decision is governed by a small number of tunable parameters. Therefore, NOD yields rapid decision-making only if it is needed by the robot’s specific task. This trait is inherent in the NOD model. Alternative approaches, such as Bayesian methods, may require additional (manual) modifications to the model on a case-by-case basis to alter its sensitivity to external inputs. In this paper, we will learn a DNN for automatic, context-aware tuning of NOD parameters.
4 Neural NOD for Interactive Robotics
In this section, we present our main contribution: Learning Neural NOD from inverse dynamic games for fast decision-making in multi-agent non-cooperative interactions.
4.1 Automatic Tuning of Game Costs using NOD
Our key insight into opinion states is that they not only model agent $i$’s attitude towards option $\ell$ (determined by the sign of ${z}^{i}_{\ell}$) but also indicate how strongly the agent prefers such an option (determined by the magnitude of ${z}^{i}_{\ell}$). Therefore, we propose to use an opinion state ${z}^{i}_{\ell}$ in a dynamic game as the coefficient of the cost term that corresponds to option $\ell$. Consequently, we may use NOD to evolve the opinion state, thus automatically tuning the game cost parameter $\theta^{i}$ in (2) such that each agent can rapidly adapt to other agents’ opinion shift as well as changes in the environment, yielding situational awareness, a crucial property for interactive robotics. To this end, we define a class of opinionated game stage costs in (2), which encodes agents’ preferences through their opinion states:
$c_{t}^{i}({x}_{t},{u}_{t};{z}^{i}_{t}):=\bar{c}_{t}^{i}({x}_{t},{u}_{t})+\sum_{\ell\in\mathcal{I}_{o^{i}}}\underbrace{{z}^{i}_{\ell,t}}_{=\theta^{i}_{\ell,t}}c_{\ell,t}^{i}({x}_{t},{u}_{t}),$ (6)
where ${z}^{i}_{\ell,t}:=\theta^{i}_{\ell,t}\in\mathbb{R}$ is the cost weight set to the opinion state of agent $i$ about opinion $\ell$, which is evolved by a state-dependent NOD model, $c_{\ell,t}^{i}({x}_{t},{u}_{t})\geq 0$ is the associated basis cost term, and $\bar{c}_{t}^{i}({x}_{t},{u}_{t})$ is the residual cost term that encodes the remaining task specifications and is independent of the options faced by the agent. We refer to a dynamic game equipped with the above cost as an opinion-guided dynamic game (ODG).
Example 1
We illustrate our technical approach with an autonomous racing example. We conduct simulated races on the 1:1 reconstructed Thunderhill Raceway located in Willows, CA, USA (see Fig. 6). In the race, both ego and rival vehicles are constrained to remain within the track boundaries (i.e., at least one wheel is inside the track limit), while only the ego vehicle has the responsibility to avoid a collision. We model the vehicle motion using the 4D kinematic bicycle model [54]. Similar to [47], ego’s basis cost terms (and associated weights) include incentivizing overtaking ($\theta^{e}_{\rm{ov}}$), following ($\theta^{e}_{\rm{fl}}$), deviating to the inside ($\theta^{e}_{\rm{in}}$) or outside ($\theta^{e}_{\rm{ot}}$); and the rival ones include incentivizing blocking the ego ($\theta^{r}_{\rm{bl}}$), and deviating to the inside ($\theta^{r}_{\rm{in}}$) or outside ($\theta^{r}_{\rm{ot}}$). This leads to a 7D joint opinion space. The residual cost terms of each player capture optimizing lap time, enforcing safety, and regulating control efforts.
Figure 2: Computation graph of the inverse game for training a Neural NOD model illustrated with the autonomous racing example.
4.2 Neural Synthesis of NOD from Inverse Dynamic Games
The key question of using NOD in an ODG (6) for physical interaction is: How should NOD parameters change adaptively in response to the evolution of physical state ${x}$? Among the few attempts to address this issue, recent work [22] captures physical information in NOD parameter design via value functions of a dynamic game, assuming that each agent faces a set of mutually exclusive options. Cathcart et al. [9] uses a manually-designed NOD model for human–robot corridor passing. In this section, we present an inverse game approach to synthesize a Neural NOD model for general ODG s with no additional assumptions on each agent’s specific task and set of options.
A discrete-time Neural NOD model (c.f. (5)) is defined as:
${z}_{t+1}$ $={g}({z}_{t};\eta_{t}),$ $\forall t>0$ (7a)
$\eta_{t}$ $=h_{\eta}({x}_{t};\phi),$ $\forall t>0$ (7b)
${z}_{0}$ $=h_{{z}_{0}}({x}_{0};\phi_{0}),$ (7c)
where $\eta_{t}=(\eta^{1}_{t},\ldots,\eta^{N_{a}}_{t},{\lambda}_{t})$, with $\eta^{i}_{t}=({d}^{{i}}_{t}$, ${b}^{i}_{t}$, $\alpha^{{i}}_{t}$, $\gamma^{{i}}_{t}$, $\beta^{{i}}_{t}$, $\delta^{i}_{t})$, is the vector that aggregates all NOD parameters. In our proposed scheme, $\eta_{t}$ is predicted by DNN $h_{\eta}$ with parameters $\phi$ and input ${x}_{t}$. Given an initial state ${x}_{0}$, the NOD model is initialized with opinion ${z}_{0}$, which is predicted by a separate DNN $h_{{z}_{0}}$ parameterized by $\phi_{0}$. The predicted ${z}_{0}$ can also be viewed as prior information of agents’ opinions, with which we can initialize NOD with a more informative opinion than a neutral one (i.e., ${z}_{0}=0$).
Remark 2 (DNN features)
The Neural NOD framework is agnostic to the specific DNN feature space or architecture. For ease of exposition, we use the current state ${x}_{t}$ as the input feature of DNN $h_{\eta}(\cdot;\phi)$. Alternatively, we may use an LSTM [17] or transformer [45] model for $h_{\eta}(\cdot;\phi)$, which would then predict NOD parameters based on a state history.
Our technical approach towards synthesizing a Neural NOD model is inspired by recent work on inverse dynamic games, but incorporates several key distinctions. Existing inverse game approaches, such as [42, 29, 33], are predominantly limited to learning a static cost parameter $\theta$ from a single gameplay trajectory data. As a result, the inverse game problem needs to be re-solved when the environment (e.g., the initial condition or the other agent’s intent) changes, presenting computational hurdles for real-time deployment. In contrast, our proposed approach involves offline training of a Neural NOD model from a dataset containing many gameplay trajectories. Subsequently, this model is deployed online in an ODG to dynamically adjust game cost weights based on evolving physical states, thereby facilitating situational awareness among the gameplay agents. The inverse game training needs not be performed online, provided that the deployment scenario falls within the generalization capability of the DNN that parameterizes the NOD. In addition, the online computation speed of agents’ policies via solving an ODG with learned Neural NOD is close to that of a standard dynamic game, as long as the DNN used by Neural NOD adopts fast inference. The Neural NOD framework also differs from the game-induced NOD (GiNOD) [22] in that GiNOD assumes mutually exclusive options, requires (potentially expensive) online optimization to synthesize NOD parameters, and is incapable of leveraging prior knowledge such as expert demonstrations. Our approach overcomes these drawbacks by learning a generic DNN-parameterized NOD model from interaction data.
We cast Neural NOD training as an inverse dynamic game. The corresponding MLE problem is formulated as:
$\max_{\phi,\phi_{0},\{({\mathbf{x}}_{n},{\mathbf{u}}_{n})\}_{n\in[N]}}{}$ $L:=\textstyle\sum_{n=1}^{N}p(\mathbf{{y}}_{n}\mid{\mathbf{x}}_{n},{\mathbf{u}}_{n})$ (8)
s.t. $({\mathbf{x}}_{n},{\mathbf{u}}_{n})\in\Gamma(\phi,\phi_{0}),\quad\forall n\in[N],$
where $L$ is the imitation objective defined by a set of (partial) observations $\{\mathbf{{y}}_{n}\}_{n\in[N]}$ and the corresponding set of state and control trajectories $\{({\mathbf{x}}_{n},{\mathbf{u}}_{n})\}_{n\in[N]}$, wherein each pair is at an equilibrium of the ODG governed by Neural NOD with parameter $(\phi,\phi_{0})$. These parameters can be learned by solving MLE (8) using standard gradient-based methods [8, 27], in which gradients $\nabla_{\phi}L$ and $\nabla_{\phi_{0}}L$ can be obtained with automatic differentiation (e.g., [40, 34]) by backpropagating through the inverse game computation graph, as shown in Figure 2. Note that MLE (8) learns DNN $h_{\eta}$ and $h_{{z}_{0}}$, which take as input arbitrary state ${x}\in\mathbb{R}^{n_{x}}$, while MLE (3) only works for a specific initial state ${x}_{0}$. Our proposed training pipeline solely requires that the ODG is formulated with any existing differentiable dynamic game frameworks, such as open-loop Nash [41], feedback Nash [29], generalized Nash [33], stochastic Nash [36, 32], open-loop Stackelberg [19], and feedback Stackelberg [30].
Example 2
We use inverse game approach [29] with a negative log-likelihood objective, which yields an approximate feedback Nash equilibrium for the ODG. We parameterize the Neural NOD as a full-connected feed-forward DNN, of which the features include all vehicles’ physical states. We implement the training pipeline with Flax [16] and train the Neural NOD with Adam [27].
4.3 Properties of Neural NOD
In this section, we provide stability analysis of Neural NOD around the neutral opinion ${z}=0$, i.e., where all agents are indecisive. In particular, we show in the following theorem that the learned self-reinforcement gains $\alpha$ and attention ${\lambda}$ can be adjusted analytically so that ${z}=0$ is an unstable equilibrium of Neural NOD, that is, indecision (i.e., deadlocks) can be broken with an arbitrarily small input from the bias ${b}$. We start with a technical lemma, which ensures that pitchfork bifurcation is possible with a Neural NOD model.
Lemma 1
Define Jacobian matrix ${\mathbf{J}}_{0}=\left.{\mathbf{J}}(\operatorname{col}(\{{S}^{i}({z}^{{i}};\eta^{i})\}_{{i}\in\mathcal{I}_{a}}))\right|_{z=0}$. If $\exists{i}\in\mathcal{I}_{a}$, $\ell\in\mathcal{I}_{o^{i}}$ such that $\alpha^{i}_{\ell}+\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})>0$, where $\bar{{\mathbf{J}}}_{0}=\left.{\mathbf{J}}(\operatorname{col}(\{\bar{{S}}^{i}({z}^{{i}};\eta^{i})\}_{{i}\in\mathcal{I}_{a}}))\right|_{z=0}$, $\bar{{S}}^{i}({z}^{{i}};\eta^{i})={S}_{1}\Big{(}\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\gamma^{{i}{j}}_{\ell}{z}^{{j}}_{{\ell}}\Big{)}+\sum_{{p}\in\mathcal{I}_{\theta_{i}}\setminus\{{\ell}\}}{S}_{2}\Big{(}\beta_{\ell{p}}^{{i}}{z}^{{i}}_{{p}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\delta^{{i}{j}}_{\ell{p}}{z}^{{j}}_{{p}}\Big{)}$, and $\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})$ is the eigenvalue of $\bar{{\mathbf{J}}}_{0}$ at the same location in the spectral matrix of $\bar{{\mathbf{J}}}_{0}$ as $\alpha^{i}_{\ell}$ in $\mathcal{A}=\operatorname{diag}\{\alpha^{i}_{\ell}\}_{i\in\mathcal{I}_{a},\ell\in\mathcal{I}_{o^{i}}}$, then ${\mathbf{J}}_{0}$ has at least one eigenvalue with a positive real part, i.e., ${}^{+}(\sigma({\mathbf{J}}_{0}))\neq\emptyset$.
Proof
Note that ${\mathbf{J}}_{0}=\mathcal{A}+\bar{{\mathbf{J}}}_{0}$. Since $\mathcal{A}$ is diagonal, each eigenvalue of ${\mathbf{J}}_{0}$ takes the form of $\sigma^{i}_{\ell}({{\mathbf{J}}}_{0})=\alpha^{i}_{\ell}+\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})$. Therefore, by choosing one $\alpha^{i}_{\ell}$ such that $\alpha^{i}_{\ell}+\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})>0$, we can ensure that ${\mathbf{J}}_{0}$ has at least one eigenvalue with positive real part. ∎
Now, we are ready to state our main theoretical result, which guarantees breaking of indecisions (deadlocks) when using a Neural NOD model for game-theoretic planning.
Theorem 4.1 (Guaranteed Indecision Breaking)
Consider the continuous-time Neural NOD model $\dot{{z}}={g}_{c}({z};\eta)$, where ${g}_{c}=\operatorname{col}({g}_{c}^{1},\ldots,{g}_{c}^{N_{a}})$ with ${g}_{c}^{i}(\cdot;\eta^{i})$ defined in (4) and $\eta^{i}$ defined in (7). If $\exists{i}\in\mathcal{I}_{a}$, $\ell\in\mathcal{I}_{o^{i}}$ such that $\alpha^{i}_{\ell}+\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})>0$, then the following results hold:
  • When ${b}=0$, if ${\lambda}>\|\operatorname{col}(\{{d}_{i}\}_{{i}\in\mathcal{I}_{a}})\|_{\infty}/\max{}^{+}(\sigma({\mathbf{J}}_{0}))$, then the neutral opinion ${z}=0$ is an unstable equilibrium of the Neural NOD. Moreover, the opinion state departs ${z}=0$ (locally) at an exponential rate with an arbitrarily small bias ${b}$,
  • When ${b}\neq 0$, the neutral opinion ${z}=0$ is not an equilibrium of the Neural NOD and the pitchfork unfolds, i.e., the model is ultrasensitive at ${z}=0$.
Proof
Linearizing Neural NOD at neutral opinion ${z}=0$ gives $\dot{{z}}=(-D+{\mathbf{J}}_{0}){z}+{b}$, where $D=\operatorname{blkdiag}(\{D^{i}\}_{i\in\mathcal{I}_{a}})$. When ${b}=0$, indecision ${z}=0$ is an equilibrium. Since ${}^{+}(\sigma({\mathbf{J}}_{0}))\neq\emptyset$, if the attention variable ${\lambda}$ is chosen such that ${\lambda}>{\lambda}^{*}=\|\operatorname{col}(\{{d}_{i}\}_{{i}\in\mathcal{I}_{a}})\|_{\infty}/\max{}^{+}(\sigma({\mathbf{J}}_{0}))$, then matrix $-D+{\mathbf{J}}_{0}$ has one eigenvalue with positive real part, i.e., indecision ${z}=0$ is exponentially unstable. Thus, an indecision-breaking pitchfork bifurcation occurs at critical attention value ${\lambda}^{*}$. When ${b}\neq 0$, the pitchfork bifurcation unfolds, as predicted by the unfolding theory [15]. ∎
4.4 Using Neural NOD for Split-Second Decision-Making
Once a Neural NOD model is learned in an offline manner from data, it can be deployed in a plug-and-play fashion in an online ODG solver for fast and flexible multi-agent decision-making Moreover, Theorem 4.1 offers a systematic approach to avoid indecisions in an ODG by modifying the self-reinforcement gains $\alpha$ and attention ${\lambda}$ as a function of Neural NOD parameters. Finally, our proposed game-theoretic decision-making framework is flexible to mix decisive and indecisive players. For example, in car racing, we would like to learn a policy for a decisive ego agent, while not assuming the rivals are equally decisive at all times. In this case, we can jointly learn an opinion dynamics model for the ego and rivals, where the ego uses a Neural NOD model for fast and decisive motion planning, and rivals use a general DNN $\tilde{g}$ to generate the next opinion state based on the current opinion and physical state, i.e., ${z}^{r}_{t+1}=\tilde{g}({z}_{t},{x}_{t})$. By removing the inductive bias of NOD fast decision-making for the rival agents, we can more effectively predict their actions when they are slow at making decisions.
5 Simulation Results
We use our proposed Neural NOD model, trained on both synthetic and human datasets, to automatically tune the game cost weights within the ILQGame planning framework [14], which computes an approximate feedback Nash equilibrium strategy [5, Ch. 6] for each agent. We evaluate its performance in simulated autonomous racing scenarios (running example). For both planning and simulation, we employ the 4D kinematic bicycle model [54], discretized with a time step of $\Delta t=0.1$ s, to describe the motion of the vehicles. All planners are implemented using JAX [34] and run in real-time at a frequency of 10Hz on a desktop with an AMD Ryzen 9 7950X CPU.
Hypotheses. We make three hypotheses that highlight the advantages of Neural NOD.
  • H1 (Performance). The Neural NOD model leads to safer and more efficient robot motion than non-NOD baselines.
  • H2 (Generalization). The Neural NOD model generalizes better to other agents’ behaviors that are out of distribution.
  • H3 (Human data compatibility). The inverse game training pipeline can effectively learn a Neural NOD model from noisy human data.
Baselines. To validate these hypotheses, we compare Neural NOD model (7) against two baseline methods. All neural networks use the same architecture and number of parameters (fully-connected, feed-forward, and 5 hidden layers with 256 neurons each).
  • Multi-layer perceptron learned from an inverse game (MLP-IG). A mapping from feature ${x}_{t}$ to cost weights $\theta_{t}$, which is parameterized by a generic multilayer perceptron (MLP) model and learned with an inverse dynamic game. This approach is briefly explored in the prior state-of-the-art inverse game approach Liu et al. [33], which learns a tiny neural network that predicts game objectives.
  • End-to-end behavior cloning (E2E-BC). An end-to-end control policy (i.e., one that takes as input state ${x}_{t}$ and returns agents’ control ${u}_{t}$) learned with behavior cloning and supervised learning [7].
Metrics. We consider the following performance metrics:
  • Safe rate (SR). A ratio defined as $N_{\text{safe}}/N_{\text{trial}}\times 100\%$, where $N_{\text{safe}}$ is the number of safe trials—those in which the ego car stays within the track limits and avoids collisions with its rival at all times—and $N_{\text{trial}}$ is the total number of trials.
  • Overtaking rate (OR). A ratio defined as $N_{\text{overtake}}/N_{\text{trial}}\times 100\%$, where $N_{\text{overtake}}$ is the number of safe trials in which the ego car successfully overtakes the rival and maintains a lead over it at the end of the trial.
  • Average end-time leading distance (AELD). The distance between the ego car and its rival at the end of a trial, measured in meters.
  • Average accumulated steering rate difference (AASD). The accumulated steering rate difference is defined as $\sum_{k=0}^{\tilde{T}-1}|\omega_{k+1}-\omega_{k}|$ for each trial, where $\tilde{T}$ is the trial horizon, and $\omega_{k}$ is the steering rate control of the ego car at time $k$. This metric assesses the overall turning smoothness of the ego car, with a higher AASD indicating a greater likelihood of spinning out.
We will report SR and OR in terms of percentage, and AELD and AASD in terms of mean and standard deviation calculated across all simulation trials.
5.1 Synthetic dataset
In the first experiment, we learn all policies from a synthetic dataset with racing demonstrations obtained by running a game-theoretic solver. The expert game solver, inspired by ILQGame [14] with customizations to account for the Frenet coordinate [25] and track limits, computes an approximate feedback Nash equilibrium solution in real-time.
Figure 3: Simulation snapshots and the time evolution of game cost weights, when the ego vehicle uses the Neural NOD learned from the synthetic dataset. Planned future motions are displayed with transparency. The racing line is plotted in dashed black. The ego car made a timely decision to speed up and move to the outside, safely overtaking the rival.
Figure 4: Simulation snapshots and the time evolution of game cost weights, when the ego vehicle uses the baseline MLP learned from the synthetic dataset. The ego made a greedy choice by driving on the inside to stay closer to the racing line and hesitated to overtake the rival when approaching. Both factors led to a collision between the ego and the rival.
We first evaluated each policy with in-distribution rival behaviors, where the rival used an ILQGame policy whose blocking cost weight was randomized within the same range ($\theta^{r}_{\rm{bl}}\in[0,30]$) as the training data. We simulated 100 races with randomized initial states and rival cost weights, and the statistics are shown in Table 1. The proposed Neural NOD outperformed both baselines in all metrics. We then simulated another 100 randomized races against a more aggressive rival, whose blocking cost weight was randomized within the range $\theta^{r}_{\rm{bl}}\in[60,80]$. As shown in Table 2, in this more challenging, out-of-distribution setting, Neural NOD led the baselines by a significant margin in terms of safe rate, overtaking rate, and AELD. It maintained the same level of smoothness as the MLP baseline, as indicated by the tiny difference in AASD. The E2E-BC baseline yielded a significantly lower safe rate, confirming the well-known generalization issue of behavior cloned policies when deployed in previously unseen scenarios. Finally, in 100 additional randomized races where the rival used the E2E-BC policy—a shift from the game-based policy in the synthetic dataset—the Neural NOD continued to outperform the baselines in all metrics. Those results jointly validated H1. The results in Table 2 and Table 3 validated H2.
Remark 3 (Information Privilege)
Although the rival’s ILQGame cost weights were not accessible to the ego’s game policy (Neural NOD and MLP-IG), the two players’ game solvers still shared a set of common knowledge including the basis cost functions and system dynamics, which can constitute privileged information of the game-based policies over the E2E-BC baseline. Therefore, we provided additional simulation results in Table 3 with the rival using the E2E-BC policy, in which case the game-based policies no longer have an information privilege.
[…]
MLP-IG $\mathbf{95\%}$ $93.68\%$ $42.87\pm 24.30$ $1.32\pm 0.33$
E2E-BC $90\%$ $88.89\%$ $39.39\pm 26.68$ $1.87\pm 0.46$
Neural NOD (ours) $\mathbf{95\%}$ $\mathbf{96.84\%}$ $\mathbf{50.85\pm 29.91}$ $\mathbf{1.29\pm 0.38}$
Table 1: Results obtained from 100 randomized races with in-distribution rival behaviors and policies trained on the synthetic dataset.
[…]
MLP-IG $82\%$ $62.20\%$ $4.50\pm 37.84$ $\mathbf{1.07\pm 0.58}$
E2E-BC $66\%$ $60.61\%$ $3.64\pm 33.71$ $2.21\pm 0.47$
Neural NOD (ours) $\mathbf{91\%}$ $\mathbf{76.92\%}$ $\mathbf{7.60\pm 42.25}$ $1.09\pm 0.66$
Table 2: Results obtained from 100 randomized races in an out-of-distribution evaluation with more aggressive rival behaviors.
[…]
MLP-IG $70\%$ $61.43\%$ $11.77\pm 35.00$ $1.48\pm 0.43$
E2E-BC $55\%$ $61.82\%$ $14.32\pm 35.07$ $2.52\pm 0.57$
Neural NOD (ours) $\mathbf{87\%}$ $\mathbf{74.71\%}$ $\mathbf{16.34\pm 52.49}$ $\mathbf{1.36\pm 0.36}$
Table 3: Results obtained from 100 randomized races with the rival using the behavior-cloned policy trained on the synthetic dataset.
[…]
MLP-IG $78\%$ $53.85\%$ $14.49\pm 42.95$ $1.23\pm 0.64$
E2E-BC $62\%$ $75.81\%$ $12.35\pm 34.77$ $2.02\pm 0.55$
Neural NOD (ours) $\mathbf{81\%}$ $\mathbf{82.72\%}$ $\mathbf{15.94\pm 34.55}$ $\mathbf{1.09\pm 0.52}$
Table 4: Results obtained from 100 randomized races with ego policy trained on the human-generated dataset. The rival uses a game policy with randomized cost weights.
In Figure 3 and Figure 4, we examine one representative simulated race with the ego using the Neural NOD and the baseline MLP, respectively. In both cases, vehicles started from the same initial conditions, and the rival used a game policy whose cost parameters were not accessible to the ego’s policy. The ego using Neural NOD rapidly formed a strong opinion in favor of overtaking the rival from the outside. This resulted in a smooth, decisive, and safe overtaking maneuver. On the other hand, the MLP model produced a different ego behavior: the intention (encoded by $\theta^{e}_{\rm{ov}}$) to overtake the rival started high but gradually decreased as the ego approached the rival and hesitated to carry out the overtaking. Additionally, in contrast to the NOD decision, the ego greedily chose to overtake from the inside—a position closer to the time-optimal racing line but more prone to collisions due to the narrow corridor between the rival and track limit. Indeed, as the rival also attempted to defend the inside to prevent the ego from overtaking, the ego was left with no space to dodge but collided with the rival.
Figure 5: Comparing a simulated gameplay trajectory against the groundtruth. Top: Groundtruth trajectories of the ego and rival. Middle: Simulation snapshots when the ego vehicle uses the Neural NOD to race against a rival, whose motion is replayed from the groundtruth data. Bottom: Time evolution of game cost weights tuned by the Neural NOD.
5.2 Human-generated dataset
Next, we learn all policies from a dataset with racing demonstrations performed by human drivers. To obtain such racing data, we leveraged a driving simulator (shown in the upper right corner of Figure 2), where the ego car was driven by human participants of various skill levels against a rival using a reactive RL-based policy in Carla [11]. For evaluation, we simulated 100 races with randomized initial states and a rival using a game policy with randomized cost weights inaccessible to the ego’s planner. The racing statistics are shown in Table 4. The Neural NOD again outperformed both baselines in all metrics. This result coincides with those obtained based on the synthetic dataset, confirming the ability of our proposed inverse game framework to learn a Neural NOD from noisy human data (H3). In Figure 5, we compare a simulation trajectory generated by Neural NOD with the groundtruth replay. Under this initial condition when the rival tried to defend from the inside (left), the Neural NOD decided to attack from the outside (right), controlling the ego car to successfully overtake the rival. Note that this decision closely resembles the groundtruth human demonstration, also displayed in Figure 5.
5.3 The full endurance race
In the last example, we test Neural NOD in an endurance race inspired by the popular video game Real Racing 3. In this race, the ego car is required to complete one full lap with dual objectives: overtaking as many rivals as possible while minimizing the lap time. Each time a rival player is overtaken, it will be eliminated from the race, and a new rival will be initialized in front of the ego car. We use DNN $h_{{z}_{0}}$ (c.f. (7c)) to reset the opinion states when the rival is respawned. Using the Neural NOD, the ego completed the race in $101.4$ s and overtook $9$ rivals without collision. With the MLP baseline, the ego finished the race in $114.2$ s, overtaking $7$ rivals but incurred $2$ crashes. Finally, under the behavior cloning baseline, the race was concluded in $110.1$ s, during which the ego overtook $9$ rivals but with $3$ crashes. This result validated H1.
Figure 6: Ego trajectory and velocity profile of the full endurance race at the Thunderhill Raceway. Transparent footprints denote the planned motion of the ego (red) and rival (silver). The black arrow and the grid indicate the track direction and finish line, respectively.
6 Limitations and Future Work
In this paper, we consider a restricted setting where Neural NOD is used to automatically tune the (continuous) cost weights of a dynamic game solver. As shown in recent work [22], an opinion state can be interpreted as a simplex (e.g., a probability vector), and thus a Neural NOD model is also capable of coordinating agents over integer-valued options such as leadership assignment in a Stackelberg setting [26, 19]. In addition, while we have demonstrated that the attention $\lambda$ governed by the black-box DNN generally performed well for car racing, we may use an explicit attention model that enables excitable decision-making by adding an extra slower negative feedback loop [28, Sec. 7.2.2]. Such enhanced flexibility in excitable decision-making allows agents to timely forget prior decisions and thus never get stuck in a decision that may no longer be safe and/or efficient as the environment has evolved rapidly. Finally, we see an open opportunity to use Neural NOD for shared autonomy, e.g., AI-assisted car racing [49, 38], in which the human and robot simultaneously provide control inputs to the system while interacting with other agents. By modeling the intents of the human and robot as opinion states and planning robot motion in the joint intent–physical space [18], we can achieve not only value alignment, i.e., the robot infers and adopts the human’s goals, but also automation transparency, i.e., the human is aware of the robot’s current intent, both rapidly and decisively with Neural NOD.
7 Conclusion
In this paper, we introduced a Neural NOD model for game-theoretic robot motion planning in split-second, and an inverse game approach to learn such a model from data. We also provided a constructive procedure to adjust Neural NOD parameters online such that breaking of indecision is guaranteed. Through extensive simulation studies of autonomous racing based on real-world circuit and human-generated interaction data, we demonstrated that a dynamic game policy guided by a Neural NOD can consistently outperform state-of-the-art imitation learning and data-driven inverse game policies. Beyond fully-automated robot interactions, the framework of Neural NOD provides an open opportunity to enable fast, decisive, and transparent value alignment for broader human-AI shared control systems.
Acknowledgements
This work is supported by Toyota Research Institute (TRI). It solely reflects the opinions and conclusions of its authors and not TRI, or any other Toyota entity. The authors thank Jingqi Li and Xinjie Liu for very helpful discussions on inverse dynamic games, and Alessio Franci for generously providing insights into nonlinear opinion dynamics.